OpenStudy (anonymous):

Auditions for a play in New York were held with 100 men and women trying out for the lead singer position. If twice the number of men less three times the number of women is equal to 20, how many women auditioned?

OpenStudy (texaschic101):

m + w = 100 -->m = 100 - w 2m - 3w = 20 now just sub 100 - w in for m and solve for w and you will have your answer

OpenStudy (anonymous):

thank you

OpenStudy (texaschic101):

no problem :)